November 1997  ICTCM X (Chicago)  5

Simple Examples

> BVPsolve( { ODE, BCa }, y(x) );

y(x) = 2 cos(x)

> BVPsolve( { ODE, BCb }, y(x) );

No particular solution satisfies all BCs

> BVPsolve( { ODE, BCc }, y(x) );

y(x) = 2 cos(x) + C2 sin(x)